Web208 and Web210 are currently not reachable. There appears to be a network issue with an upstream provider. We are investigating urgently.
Update 2014-01-24 08:07 UTC: The network issue has been resolved and the servers are now accessible again.
Update 2014-01-24 09:01 UTC: The servers are unreachable again. We are working with our upstream provider to restore connectivity.
Update 2014-01-24 09:12 UTC: Connectivity to both servers has been restored. Our upstream provider discovered a problem with the network switch.
An emergency maintenance has been scheduled to take place at 18:00 UTC. The network engineers at the data center will be replacing the faulty switch.
2014-01-25 1838 UTC The data center has confirmed the network switch was replaced and the machines are back online.
Web75 is currently unresponsive. We are looking into this as a matter of urgency.
2014-01-05 06:41 UTC: The server is experiencing disk problems, and we are migrating it to new hardware. The main IP address will change from 184.108.40.206 to 220.127.116.11.
Update 3pm UTC: We’re still working on migrating all data to the new hardware.
Update 5pm UTC: All MySQL and PostgreSQL databases have now been migrated to the new server. All accounts up to the letter “i” have been migrated to the new hardware and their websites have been re-enabled. We’re working on the rest of the accounts.
Update 9.45pm UTC: The home directory migration is at the letter “s”. We’re working on the rest of the accounts.
Update 3am UTC: All data has now been migrated to the new server. We will leave a proxy running on the old server for 10 days to redirect all HTTP and HTTPS traffic to the new server. If you experience any problems with your sites open a ticket and we’ll look into it asap.
We have found numerous file system errors in the system’s log that seem to be causing performance issues, we’re now taking this machine offline to perform a file system check. We’ll update this post when we have more info.
2013-09-14 19:52 UTC: The file system check has finished its first pass and we are now verifying that the remainder of the file systems are clean.
2013-09-14 20:02 UTC: The server is now back online and functioning normally.